### v1

# Joint multi-band deconvolution for Euclid and Vera C. Rubin images

11institutetext: Laboratory of Astrophysics, Ecole Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, CH-1290 Versoix, Switzerland. 11email: utsav.akhaury@epfl.ch 22institutetext: GEPI, Observatoire de Paris, Université PSL, CNRS, 5 Place Jules Janssen, 92190 Meudon, France. 33institutetext: ICC-UB Institut de Ciències del Cosmos, Universitat de Barcelona, Martí Franquès, 1, E-08028 Barcelona, Spain. 44institutetext: ICREA, Pg. Lluís Companys 23, Barcelona, E-08010, Spain. 55institutetext: Université Paris-Saclay, Université Paris Cité, CEA, CNRS, AIM, 91191, Gif-sur-Yvette, France 66institutetext: Institutes of Computer Science and Astrophysics, Foundation for Research and Technology Hellas (FORTH), Greece
Joint multi-band deconvolution for Euclid and Vera C. Rubin images
U. Akhaury 11    P. Jablonka 1122    F. Courbin 113344    J.-L. Starck 5566
(Received September 15, 1996; accepted March 16, 1997)
With the advent of surveys like Euclid and Vera C. Rubin, astrophysicists will have access to both deep, high-resolution images, and multi-band images. However, these two conditions are not simultaneously available in any single dataset. It is therefore vital to devise image deconvolution algorithms that exploit the best of the two worlds and that can jointly analyze datasets spanning a range of resolutions and wavelengths. In this work, we introduce a novel multi-band deconvolution technique aimed at improving the resolution of ground-based astronomical images by leveraging higher-resolution space-based observations. The method capitalizes on the fortunate fact that the Vera C. Rubin $r$-, $i$-, and $z$-bands lie within the Euclid $VIS$ band. The algorithm jointly deconvolves all the data to turn the $r$-, $i$-, and $z$-band Vera C. Rubin images to the resolution of Euclid by enabling us to leverage the correlations between the different bands. We also investigate the performance of deep learning-based denoising with DRUNet to further improve the results. We illustrate the effectiveness of our method in terms of resolution and morphology recovery, flux preservation, and generalization to different noise levels. This approach extends beyond the specific Euclid-Rubin combination, offering a versatile solution to improve the resolution of ground-based images in multiple photometric bands by jointly using any space-based images with overlapping filters.
Key Words.:
Deconvolution – EuclidVera C. Rubin – HST
1 Introduction
High spatial resolution, high signal-to-noise ratio, and broad wavelength coverage are all essential to most observational challenges in astrophysics. However, it is difficult, or even impossible, to make these three miracles happen simultaneously. Space telescopes, although free of atmospheric turbulence, are limited in size. Ground-based telescopes can deliver high signal-to-noise data but are affected by atmospheric turbulence and have a higher sky background. In addition, blurring by the instrumental or atmospheric Point Spread Function (PSF) differs in each type of data and varies from band to band. To capitalise on the strengths of all types of telescopes and data, it is crucial to develop robust deconvolution techniques that can remove blurring by the PSF and that can optimize the signal-to-noise ratio of the final reconstruction, by combining all observations accounting for the different bands and PSFs.
Due to the presence of noise, image deconvolution is a challenging ill-posed inverse problem that requires regularization for a well-defined solution. Early approaches in the field acknowledged this issue, proposing solutions such as minimising the Tikhonov function (Tikhonov & Arsenin, 1977) or maximising the entropy of the solution (Skilling & Bryan, 1984). Bayesian methods also emerged, including the Richardson-Lucy algorithm applied to early Hubble Space Telescope (HST) data (Richardson, 1972; Lucy, 1974). A novel approach proposed by Magain et al. (MCS; 1998) separated point sources from extended ones and used a narrow PSF for deconvolution to achieve an improved resolution suitable for the chosen pixel sampling. This approach was enhanced with wavelet regularization for the extended channel (Cantale et al., 2016), and further refined by Michalewicz et al. (STARRED; 2023), who employed Starlets, an isotropic wavelet basis (Starck et al., 2015), to regularize the solution. There have also been efforts to jointly deconvolve multiple astronomical observations of the same sky region (Donath et al., 2023). Furthermore, Ingaramo et al. (2014) explored the combination of multiple sources by demonstrating the application of Richardson-Lucy deconvolution to merge high-resolution, high-noise images with low-resolution, low-noise images.
A notable advancement in astronomical deconvolution was the use of deep learning. Once trained, deep learning-based methods also offer significant computational efficiency compared to traditional approaches. U-Nets (Ronneberger et al., 2015) have gained popularity for their nonlinear processing capabilities and multi-scale architecture. Expanding on U-Nets, Sureau et al. (2020) introduced the Tikhonet method for deconvolving optical galaxy images, demonstrating superior performance over sparse regularization methods in terms of mean squared error (MSE) and a shape criterion that assesses galaxy ellipticity. Nammour et al. (2022) enhanced Tikhonet by incorporating a shape constraint into the loss function, achieving further improvements. Another powerful architecture, Learnlet (Ramzi et al., 2023), combines the strengths of wavelets and U-Nets while offering a fully interpretable neural network with minimal hallucination. In our previous work (Akhaury et al., 2022, 2024), we proposed a two-step deconvolution framework and investigated the performance of CNN and transformer-based denoisers. We concluded that a Swin transformer-based U-Net (SUNet, Fan et al. (2022)) outperformed a CNN-based U-Net (Ronneberger et al., 2015) in terms of normalised mean sqaured error(NMSE) and structural similarity index measure (SSIM).
While deconvolution is primarily used to reconstruct galaxy images at high spatial resolution in each photometric band independently, there are scenarios, particularly at low signal-to-noise ratios, where joint multi-band deconvolution can enhance the detection and characterization of systems. One such potential application is the joint multi-band deconvolution of Euclid and Vera C. Rubin images. The Vera C. Rubin Observatory is set to deliver a dataset of $500$ petabytes across multiple optical frequency bands, while Euclid will observe images spanning the optical and infrared spectrum. Interestingly, the Euclid $VIS$-band (central frequency $=715$ nm) overlaps with three of the Vera C. Rubin filters: $r$, $i$, $z$. As a space-based satellite, Euclid will produce images with sharper details due to its narrower PSF compared to Vera C. Rubin. A related study by Joseph et al. (2021) also involves jointly modelling simulated images that model observations from both Euclid and Vera C. Rubin.
In this work, we present a novel multi-band deconvolution technique designed to enhance the resolution of ground-based astronomical images by leveraging higher-resolution space-based observations. Our approach, which focusses on the joint deconvolution of Vera C. Rubin and Euclid images, effectively exploits the overlapping spectral coverage of the Rubin $r$-, $i$-, and $z$-bands with the Euclid $VIS$-band. By utilizing the Euclid $VIS$-band image as a term that provides additional information, our technique ensures that the deconvolved Rubin images retain high spatial resolution and accurate photometric measurements. The integration of deep learning-based denoising further enhances the quality of the deconvolved outputs, reducing background noise without altering the main structures of the galaxies. We generate the realistic Euclid and Vera C. Rubin simulations from HST cutouts of varying magnitudes extracted from the GOODS-N and GOODS-S surveys (Retzlaff, J. et al., 2010). The simulated Euclid-like $VIS$-band PSF was obtained from Liaudat et al. (2022), and the simulated Rubin-like $r$-, $i$-, and $z$-band PSFs were obtained from Abolfathi et al. (2021). Our method demonstrates effectiveness in terms of resolution recovery, flux preservation, and generalization across different noise levels. Through our joint deconvolution approach, we achieve resolution recovery in Vera C. Rubin simulated images close to that of HST, a feat nearly impossible with independent deconvolution of each photometric band. The potential applications of our method go beyond the Euclid-Rubin pair, providing a flexible solution to enhance the resolution of ground-based images in multiple photometric bands with any overlapping space-based filter band. This versatility is especially important as large-scale astronomical surveys gather increasing amounts of data, creating a need for effective and reliable deconvolution techniques.
In Section 2, we describe the deconvolution problem and introduce our proposed solution. The methodology for generating our dataset is detailed in Section 3. We then present the outcomes of our deconvolution algorithm in Section 4. Finally, Section 5 presents our conclusions. To support reproducible research, the codes utilised in this article are publicly accessible in Section 6.
2 The deconvolution problem
2.1 The forward model
For the three Vera C. Rubin filters, let $\mathbf{y}_{r},\mathbf{y}_{i},\mathbf{y}_{z}\in\mathbb{R}^{n\times n}$ be the corresponding observed images and $\mathbf{h}_{r},\mathbf{h}_{i},\mathbf{h}_{z}\in\mathbb{R}^{n\times n}$ be the PSFs. If $\mathbf{x}_{r}^{t},\mathbf{x}_{i}^{t},\mathbf{x}_{z}^{t}\in\mathbb{R}^{n\times n}$ denote the corresponding target images, $\ast$ denotes the convolution operation, and $\eta_{r},\eta_{i},\eta_{z}\in\mathbb{R}^{n\times n}$ denote additive noise, the observed Rubin images can then be modelled as:
$\mathbf{y}_{r}$ $=\mathbf{h}_{r}\ast\mathbf{x}_{r}^{t}+\eta_{r}$ (1)
$\mathbf{y}_{i}$ $=\mathbf{h}_{i}\ast\mathbf{x}_{i}^{t}+\eta_{i}$ (2)
$\mathbf{y}_{z}$ $=\mathbf{h}_{z}\ast\mathbf{x}_{z}^{t}+\eta_{z}$ (3)
As for Euclid, let $\mathbf{y}_{euc}\in\mathbb{R}^{n\times n}$ be the observed image, $\mathbf{x}^{t}_{euc}\in\mathbb{R}^{n\times n}$ be the target image, and $\mathbf{h}_{euc}\in\mathbb{R}^{n\times n}$ be the PSF. If $\eta_{euc}\in\mathbb{R}^{n\times n}$ denotes additive noise and $\alpha_{r},\alpha_{i},\alpha_{z}\in\mathbb{R}^{n}$ denote the corresponding fractional flux contribution from each Vera C. Rubin filter, the target and the observed images can be modelled as:
$\mathbf{x}^{t}_{euc}$ $=\alpha_{r}\mathbf{x}^{t}_{r}+\alpha_{i}\mathbf{x}^{t}_{i}+\alpha_{z}\mathbf{x}^{t}_{z}$ (4)
$\mathbf{y}_{euc}$ $=\mathbf{h}_{euc}\ast\mathbf{x}_{euc}^{t}+\eta_{euc}$ (5)
The motivation behind taking the weighted sum of the Vera C. Rubin images to model the Euclid image can be seen in Figure 1, which shows the overlap between the Rubin and Euclid filters.
Figure 1: The filter curves for Euclid and Vera C. Rubin, which display the relative filter transmission as a function of the wavelength. The Euclid $VIS$-band overlaps with three of the Vera C. Rubin filters: $r$, $i$, $z$.
2.2 The proposed solution
We formulated the following loss functions and minimize them using gradient descent to recover the optimal solutions:
$L_{r}(\mathbf{x}_{r})$ $=\frac{1}{2}\left\|\frac{\mathbf{h}_{r}\ast\mathbf{x}_{r}-\mathbf{y}_{r}}{\sigma_{r}}\right\|_{F}^{2}+\lambda_{r}\textbf{K}$ (6)
$L_{i}(\mathbf{x}_{i})$ $=\frac{1}{2}\left\|\frac{\mathbf{h}_{i}\ast\mathbf{x}_{i}-\mathbf{y}_{i}}{\sigma_{i}}\right\|_{F}^{2}+\lambda_{i}\textbf{K}$ (7)
$L_{z}(\mathbf{x}_{z})$ $=\frac{1}{2}\left\|\frac{\mathbf{h}_{z}\ast\mathbf{x}_{z}-\mathbf{y}_{z}}{\sigma_{z}}\right\|_{F}^{2}+\lambda_{z}\textbf{K}$ (8)
where $\text{where }\textbf{K}=\left\|\frac{\hskip 2.0pt\mathbf{h}_{euc}\ast\sum\limits_{c\in\{r,i,z\}}\alpha_{c}\mathbf{x}_{c}-\mathbf{y}_{euc}}{\sigma_{euc}}\right\|_{F}^{2}$ (9)
The first terms in equations 6-8 represent the data fidelity terms for each respective band, with $\sigma_{r}$, $\sigma_{i}$, $\sigma_{z}$ being the corresponding noisemaps. The second terms are the constraining terms that enforce the condition that the sum of images in the Rubin $r$-, $i$-, and $z$-band images equals the $Euclid$ $VIS$-band images, as expected from Figure 1 after flux calibration. Within these constraining terms, the individual images $x_{r}$, $x_{i}$, and $x_{z}$ are multiplied by their respective fractional flux contributions $\alpha_{r}$, $\alpha_{i}$, and $\alpha_{z}$, which represent the fractional area overlaps between their corresponding filter curves. The values of $\alpha_{r}$, $\alpha_{i}$, and $\alpha_{z}$ are obtained by integrating the area under the curves in Figure 1 and normalising them to sum up to one. The resulting values are $\alpha_{r}=0.3785$, $\alpha_{i}=0.3468$, and $\alpha_{z}=0.2746$. The denominator, $\sigma_{euc}$, denotes the $Euclid$ image noisemap. The choice of the multiplicative hyperparameters $\lambda_{r}$, $\lambda_{i}$, and $\lambda_{z}$ is described in Section 2.3.
2.3 Hyper-parameter tuning
The hyper-parameters $\lambda_{r}$, $\lambda_{i}$, and $\lambda_{z}$ are determined by varying the ratios between the second and first terms in equations 6-8. These ratios indicate the contribution coming from the constraining term. We varied the ratios from 0 (no contribution from the constraining term) to 1 (equal contribution from the constraining term) in steps of 0.01. From this experiment, the optimal solution yields the lowest MSE with a ratio of 0.3 for all three photometric bands. Subsequently, the values of $\lambda_{r}$, $\lambda_{i}$, and $\lambda_{z}$ are computed from these ratios by dividing them by K (equation 9).
2.4 Optimization
The aim is to find optimal solutions $\mathbf{\hat{x}}_{r}$, $\mathbf{\hat{x}}_{i}$, $\mathbf{\hat{x}}_{z}$ that minimise the individual loss functions:
$\mathbf{\hat{x}}_{\{r,i,z\}}=\underset{\mathbf{x}_{\{r,i,z\}}}{\operatorname{argmin}}\hskip 2.0ptL_{\{r,i,z\}}(\mathbf{x}_{\{r,i,z\}})$
which is done in an alternative and iterative manner using gradient descent:
$\mathbf{x}_{\{r,i,z\}}^{[k+1]}=\mathbf{x}_{\{r,i,z\}}^{[k]}-\beta_{\{r,i,z\}}\nabla L_{\{r,i,z\}}(\mathbf{x}_{\{r,i,z\}}^{[k]})$ (10)
where $\mathbf{x}^{[k]}$ denotes the variable $\mathbf{x}$ at $k^{th}$ iteration, and $\beta_{r},\beta_{i},\beta_{z}\in\mathbb{R}^{n}$ are the step-sizes chosen such that convergence is guaranteed (described in more detail in Section 2.5). While computing equation 10 for one band, it is assumed that the other two bands are known and remain constant. The gradients of the loss functions 6-8 are given by:
$\nabla L_{r}(\mathbf{x}_{r})$ $=\frac{\mathbf{h}_{r}^{\top}\ast(\mathbf{h}_{r}\ast\mathbf{x}_{r}-\mathbf{y}_{r})}{\left\|\mathbf{\sigma}_{r}\right\|_{F}^{2}}+\lambda_{r}\alpha_{r}\mathbf{K}_{grad}$ (11)
$\nabla L_{i}(\mathbf{x}_{i})$ $=\frac{\mathbf{h}_{i}^{\top}\ast(\mathbf{h}_{i}\ast\mathbf{x}_{i}-\mathbf{y}_{i})}{\left\|\mathbf{\sigma}_{i}\right\|_{F}^{2}}+\lambda_{i}\alpha_{i}\mathbf{K}_{grad}$ (12)
$\nabla L_{z}(\mathbf{x}_{z})$ $=\frac{\mathbf{h}_{z}^{\top}\ast(\mathbf{h}_{z}\ast\mathbf{x}_{z}-\mathbf{y}_{z})}{\left\|\mathbf{\sigma}_{z}\right\|_{F}^{2}}+\lambda_{z}\alpha_{z}\mathbf{K}_{grad}$ (13)
$\text{where }\textbf{K}_{grad}=\frac{2\mathbf{h}_{euc}^{\top}}{\left\|\mathbf{\sigma}_{euc}\right\|_{F}^{2}}\ast\left(\hskip 2.0pt\mathbf{h}_{euc}\ast\sum\limits_{c\in\{r,i,z\}}\alpha_{c}\mathbf{x}_{c}-\mathbf{y}_{euc}\right)$ (14)
2.5 Gradient descent step-size
Suppose a function $f:\mathbb{R}^{n\times n}\xrightarrow{}\mathbb{R}^{n\times n}$ which is convex and differentiable. Its gradient is Lipschitz continuous if there exists some constant C such that:
$\left\|\nabla f(\mathbf{x^{\prime}})-\nabla f(\mathbf{x})\right\|$ $\leq C\left\|\mathbf{x^{\prime}}-\mathbf{x}\right\|$
Since the loss functions 6-8 are convex and differentiable, one could find Lipschitz constants $C_{\{r,i,z\}}$ such that:
$\left\|\nabla L_{\{r,i,z\}}(\mathbf{x^{\prime}}_{\{r,i,z\}})-\nabla L_{\{r,i,z\}}(\mathbf{x}_{\{r,i,z\}})\right\|\leq C_{\{r,i,z\}}\left\|\mathbf{x^{\prime}}_{\{r,i,z\}}-\mathbf{x}_{\{r,i,z\}}\right\|$
$C_{\{r,i,z\}}\geq\frac{\mathbf{h}^{\top}_{\{r,i,z\}}\ast\mathbf{h}_{\{r,i,z\}}}{\left\|\mathbf{\sigma}_{\{r,i,z\}}\right\|_{F}^{2}}+2\lambda_{\{r,i,z\}}\alpha^{2}_{\{r,i,z\}}\mathbf{h}_{euc}^{\top}\ast\mathbf{h}_{euc}$ (15)
Once that is found, the optimal constraints on the individual step-sizes $\beta_{\{r,i,z\}}$ that ensure convergence are as follows:
$\beta_{\{r,i,z\}}$ $\leq\frac{1}{C_{\{r,i,z\}}}$ (16)
From equations 15 and 16, it is important to note that the step-size also depends on the $Rubin$ and $Euclid$ PSFs (detailed in Sections 3.2 and 3.3), the values of $\lambda_{\{r,i,z\}}$ (described in Section 2.3) and $\alpha_{\{r,i,z\}}$ (shown in Section 2.2).
3 Dataset generation
3.1 Ground truth images
We extracted HST cutout windows of dimensions $128\times 128$ pixels from GOODS-N and GOODS-S (Retzlaff, J. et al., 2010) in the $F606W$, $F775W$, and $F850LP$ bands by centering them at the centroid of the object. These HST bands were selected because their central wavelengths align with those of the $Rubin$ $r$-, $i$-, and $z$-bands, and these HST images were subsequently used to simulate the Rubin images, as explained in Section 3.2. The mosaiced HST ACS images along with the catalogue can be found at this link [^1] https://archive.stsci.edu/prepds/goods/. These HST cutouts are at a pixel scale of $0.05\arcsec$. We aimed to perform the experiments on galaxies with sizes large enough to effectively assess the impact of deconvolution on our ability to resolve their structural and morphological features, such as arms, bars, and clumps. To ensure the selection of large galaxies and exclude point-sized objects, we applied the following filtering criteria to the $F775W$ band catalogue:
  • $18<$ MAG_AUTO $<23$ (AB magnitude in SExtractor “AUTO” aperture)
[…]
  • FWHM $>10$ (full width at half maximum in pixels)
We then visually inspected and selected 92 objects that exhibited extended and complex galaxy structures. The histogram of the HST $F775W$ band magnitude for all galaxies in our dataset is shown in Figure 2. It is important to highlight that according to our simulations, the HST $F775W$ band corresponds to the Rubin $i$-band, and we will refer to it as the Rubin $i$-band throughout the text.
Figure 2: Histogram of the HST $F775W$ band magnitude for all galaxies in our dataset after filtering. Note that the HST $F775W$ band matches with the Rubin $i$-band.
3.2 Vera C. Rubin images
The simulated Vera C. Rubin-like PSFs were obtained from the second data challenge (DC2) of the Legacy Survey of Space and Time (LSST) Dark Energy Science Collaboration (DESC). The atmospheric and optical effects, as well as sensor-induced electrostatic effects, are simulated using physically motivated models, with a final adjustment to the PSF sizes to match the expected data from the Rubin Observatory, as described in detail in Abolfathi et al. (2021).
To generate the Vera C. Rubin simulated images, we first convolved the HST images in the $F606W$, $F775W$, and $F850LP$ bands with the Rubin $r$-, $i$-, and $z$-band PSFs, respectively, such that the simulated images are at the expected Rubin resolution, with a pixel scale of $0.2\arcsec$. Subsequently, we added white Gaussian noise such that our Rubin-simulated images have a signal-to-noise (S/N) ratio ranging between $12$ and $28$, with a median around 20. Based on the survey parameters outlined in Željko Ivezić et al. (2019), our simulations suggest an S/N range that corresponds to the initial few visits of the telescope. This indicates that our method could be effectively applied as soon as the first images start arriving.
3.3 Euclid images
The simulated Euclid-like $VIS$-band PSF was obtained using the WaveDiff model proposed by Liaudat et al. (2022), which changes the data-driven PSF modelling space from the pixels to the wavefront by adding a differentiable optical forward model in the modelling framework. WaveDiff outputs an approximation of the true Euclid PSF, which was derived before the actual launch of the satellite.
Next, we calculated the fractional flux contributions $\alpha_{r},\alpha_{i},\alpha_{z}$ by integrating the area under the curves in Figure 1 and normalising them to sum to 1. The resulting values, as also mentioned in Section 2.2, are $\alpha_{r}=0.3785$, $\alpha_{i}=0.3468$, and $\alpha_{z}=0.2746$. To generate the Euclid simulated images, we multiplied $\alpha_{r},\alpha_{i},\alpha_{z}$ with the HST images in the $F606W$, $F775W$, and $F850LP$ bands, respectively. The result is then convolved with the simulated Euclid PSF, such that the simulated images are at the expected Euclid resolution with a pixel scale of $0.1\arcsec$. Finally, white Gaussian noise is added such that the Euclid-simulated images have a signal-to-noise (S/N) ratio ranging between $20$ and $45$, with a median around 35. Based on the calculations presented by Euclid Collaboration et al. (2022) which assess the S/N statistics for Euclid, our simulations are conservative, implying that our method would perform well when applied to real images with higher S/N.
4 Results
The algorithm simultaneously processed the noisy simulations from the three Vera C. Rubin bands and the Euclid $VIS$-band, along with their respective PSFs. These noisy images served as initializations or first guesses for the algorithm. Subsequently, the algorithm iteratively minimised the loss functions 6-8, as detailed in Section 2.4. The algorithm was run for 200 iterations, with convergence typically observed within $50$-$100$ iterations for all images in our dataset. Figure 3 shows the convergence plot of the loss function for the deconvolved output in Figure LABEL:subfig:mcdec1.
Figure 3: The loss function for the galaxy shown in Figure LABEL:subfig:mcdec1. Convergence is guaranteed at around 100 iterations when the relative change in loss value $<10^{-3}$ and the curve is flat.
4.1 Flux leakage test
As a validation to ensure no flux leakage between channels during the joint deconvolution of Vera C. Rubin and Euclid images, we conducted a unit test. We assumed three distinct Gaussians placed separately in the Rubin channels, with the Euclid simulation being a weighted sum of these three images, resulting in three disjoint Gaussians. Post-deconvolution, the Gaussians remained intact without any structure extending beyond their boundaries. This confirms that the structure present in the deconvolved image within each Rubin band is independent of the structures in other bands and each image accurately retains only the information relevant to its specific band. Figure 4 provides a visual demonstration of these findings.
Moreover, to further validate that our method is effective for objects with non-flat spectral energy distributions (SEDs), we analyzed the transfer of information across different bands for such objects. The results are presented in Appendix A.
Figure 4: Unit test to verify that there is no leakage of flux from one channel to another. The recovered Gaussians remain at their original centers.
4.2 Deconvolved outputs
We present two examples of deconvolved images in Figure LABEL:fig:deconv_outputs, illustrating the algorithm’s capability to recover features that were lost in the original Vera C. Rubin simulations. Visually, the deconvolved outputs seem to capture the variations between different bands. Qualitatively, these outputs exhibit high quality with minimal background noise and result in clean residuals. However, in the next Section, 4.3, we demonstrate that employing a deep learning-based denoiser further enhances the already impressive results achieved initially. Table 2 presents the normalized mean squared error (NMSE) computed with respect to the ground-truth HST image for the pre-denoised and post-deonised images.
4.3 Deep learning-based denoising
After obtaining the deconvolved outputs, we feed them to DRUNet, a neural network proposed by Zhang et al. (2022) that combines U-Net (Ronneberger et al., 2015) and ResNet (He et al., 2016). U-Net is renowned for its efficiency in image-to-image translation, while ResNet excels in increasing modelling capacity through stacked residual blocks. Inspired by FFDNet (Zhang et al., 2018a), which incorporates a noise level map as input, DRUNet enhances U-Net by integrating residual blocks to improve prior denoising modelling. Similar approaches that combine U-Net and ResNet can be found in other studies (Zhang et al., 2018b; VenkateshG et al., 2018). The backbone of DRUNet is a U-Net architecture with four scales, and the schematic proposed by Zhang et al. (2022) is depicted in Figure 5.
We chose DRUNet because it is a non-blind denoiser, meaning it takes the noise map as input. This ensures that any unknown noise can be estimated and given as input for denoising. We estimate the noise level map in our deconvolved images using the scikit-image package (van der Walt et al., 2014) by calculating the average noise level within four $16\times 16$ pixel squares placed at each corner of the image. This approach ensures that only background noise is measured, avoiding any contribution from the signal. This map is then fed to the pre-trained DRUNet, along with the deconvolved image. The denoised outputs are illustrated in Figure LABEL:fig:denoise_outputs. It is observed that the denoiser exclusively eliminates noise from the background without affecting the main structure of the galaxies. Although the enhancement in image quality is marginal (since the original deconvolved image is already of high quality), the normalized mean squared error (NMSE) computed with respect to the ground-truth HST image decreases, as shown in Table 2. The most notable improvement is observed in the $z$-band images. Even though DRUNet was trained on a combination of images from BSD (Chen & Pock, 2017), Waterloo Exploration Database (Ma et al., 2017), DIV2K (Agustsson & Timofte, 2017), and Flick2K (Lim et al., 2017), it is remarkable that it works well on astronomical data, showing great generalization. Finally, the fractional error in the output flux as a function of the $i$-band magnitude is shown in Figure 6, which indicates that the mean flux error is less than $5\%$ for the entire magnitude range for all the bands.
Figure 5: The DRUNet architecture, which incorporates an additional noise level map as input and integrates U-Net (Ronneberger et al., 2015) with ResNet (He et al., 2016).”SConv” stands for strided convolution and ”TConv” stands for transposed convolution. Image credits: Zhang et al. (2022).
Figure 6: Fractional error in the output flux as a function of the $i$-band magnitude (which is chosen in order to have the same scale on the x-axes). The dots correspond to the individual galaxies, and the grey line is the best-fit line after binning the magnitude values.
Table 1: NMSE with respect to HST images.
NMSE $r$-band $i$-band $z$-band
Pre-denoising 0.059 0.041 0.053
Post-denoising 0.058 0.038 0.038
% improvement 1.69% 7.32% 28.3%
 [^2] NMSE is calculated for the pre-denoised and post-denoised images with respect to the ground-truth HST image.
4.4 Improvement compared to independent deconvolution of each band
To demonstrate the advantages of jointly deconvolving multiple photometric bands, we performed independent deconvolution of each band using the deconvolution framework introduced in Akhaury et al. (2024). The method proceeds in two steps: it first deconvolves the input using Tikhonov regularization (Tikhonov & Arsenin, 1977), and denoises the result using SUNet (Fan et al., 2022), a state-of-the-art Swin transformer-based architecture. The results, presented in Figure 8, show that the joint deconvolution method outperforms the independent deconvolution of individual bands. The joint method enables us to leverage the correlation between the different bands and the space-based image, thus improving the final output.
5 Conclusion
In this work, we have presented a novel image deconvolution technique designed to improve the resolution of multi-band ground-based data by leveraging higher-resolution space-based observations. Our approach, which focusses on the joint deconvolution of Vera C. Rubin and Euclid images, effectively exploits the overlapping spectral coverage of the Rubin $r$-, $i$-, and $z$-bands with the Euclid $VIS$-band. Through rigorous testing, we demonstrated that our iterative algorithm successfully recovers fine details while preserving the flux for each band. Different noise levels were tested, and the resolution achieved by ground-based data is close to that of HST. Our results indicate that joint deconvolution of all data outperforms independent deconvolution of individual photometric bands using existing state-of-the-art methods. By utilising the Euclid $VIS$-band image as a term that provides additional information, our technique ensures that the deconvolved Rubin images retain high spatial resolution and accurate photometric measurements. The integration of deep learning-based denoising using DRUNet enhances the quality of the deconvolved output, further reducing background noise without altering the main structures of the galaxies.
The potential applications of our method extend beyond the Euclid-Rubin pair, offering a versatile solution to improve the resolution of ground-based images in multiple photometric bands as long as there exists a space-based image of the same field of view in a band that encompasses all ground-based filters. In the future, we intend to test our deconvolution method on images of the Perseus cluster by using ground-based observations from the CFHT (Canada–France–Hawaii Telescope) and space-based observations from the Euclid ERO (Early Release Observations) public release.
6 Data availability
For the sake of reproducible research, the codes used for this article are publicly available online.
  1. 1.
    The ready-to-use version of our joint deconvolution method [^3] https://github.com/utsav-akhaury/Multiband-Deconvolution/tree/main.
  2. 2.
    The DRUNet repository from Zhang et al. (2022)  [^4] https://github.com/cszn/DPIR.
Acknowledgements.
This work was funded by the Swiss National Science Foundation (SNSF) under the Sinergia grant number CRSII5_198674. This work was supported by the TITAN ERA Chair project (contract no. 101086741) within the Horizon Europe Framework Program of the European Commission, and the Agence Nationale de la Recherche (ANR-22-CE31-0014-01 TOSCA).
Figure 8: The galaxy in Figure LABEL:subfig:mcdec2 deconvolved using two different approaches, illustrating that joint deconvolution outperforms independent deconvolution of individual photometric bands. The joint method allows us to leverage the correlation between the different bands and the space-based image, thus improving the final output. The first column displays the Euclid $VIS$ image, followed by Vera C. Rubin simulations in the $r$, $i$, and $z$ bands in the second column. The independently deconvolved SUNet outputs for the three bands are shown in the third column. The fourth column shows the corresponding joint deconvolution outputs followed by denoising with DRUNet. The ground-truth HST images are shown in the fifth column.
Appendix A Generalization to objects with non-flat spectral energy distribution (SED)
To assess the impact of information transfer across different bands for objects with a non-flat spectral energy distribution (SED), we conducted an experiment where we successively replaced the Rubin bands with pure noise, ensuring that no galaxy signal was present. We then analyzed how features from the high-resolution Euclid image propagated into the deconvolved Rubin bands. The outputs are shown in Figure LABEL:fig:noise_test.
Our results confirm that reconstructed features appear only in bands where the original galaxy signal is present. This demonstrates that the algorithm does not artificially imprint Euclid information onto Rubin bands lacking real data. In cases where a structure is visible in the Euclid image but absent from one or more Rubin bands, the problem becomes degenerate: the feature could either be entirely attributed to a single band or distributed across multiple bands. However, our findings indicate that the outputs are directly influenced by the input data in each band rather than being dictated solely by the high-resolution Euclid image. Our results demonstrate that the joint deconvolution method effectively utilises the available signal in each band while respecting the constraints imposed by the data. This also confirms that the method would work for objects with non-flat SEDs, where signal may be present in only one band but absent in others, ensuring that features are accurately transferred according to their actual distribution across the bands.
As mentioned in Section 4, we found that the algorithm converges within $200$ iterations when the signal is available in all three bands. When one band is replaced with a noisemap, as shown in Figure LABEL:subfig:noise_test1, convergence requires approximately $1000$ iterations. Replacing a second band with a noisemap, as seen in Figure LABEL:subfig:noise_test2, further increases the iteration count to around $5000$. This demonstrates that incorporating more data across different bands significantly accelerates the convergence of the loss functions as the algorithm can benefit by capturing the correlations between these bands.
Appendix B Plug-and-Play ADMM
Plug-and-Play Alternating Direction Method of Multipliers (PnP ADMM) has emerged as a powerful framework for solving inverse problems by combining iterative optimization techniques with deep learning-based priors. Originally developed for convex optimization problems with linear equality constraints (Boyd et al. 2010), ADMM decomposes the minimization process into sequential sub-problems, typically involving a data fidelity term and a regularization term, followed by an update of the dual variable. Previous works (Venkatakrishnan et al. 2013; Sreehari et al. 2016; H. Chan et al. 2016) have interpreted these sub-steps as an inversion step followed by a denoising step, coupled via the augmented Lagrangian term and the dual variable. The ADMM PnP approach extends this idea by replacing the proximal operator related to the prior with a deep neural network (DNN) trained as a denoiser (Meinhardt et al. 2017; Bigdeli et al. 2017; Gupta et al. 2018; Sureau et al. 2020), allowing for greater flexibility in handling complex image priors. Compared to direct deep-learning-based inverse models, PnP ADMM offers several advantages: (1) it decouples the inversion step from the DNN, enabling the inclusion of additional convex constraints that can be efficiently handled via optimization, (2) it reduces the cost of learning by focusing solely on training a denoiser rather than multiple networks, as seen in unfolding approaches, and (3) by iterating between denoising and inversion, it ensures that the network output remains consistent with the observed data. In this work, we integrate the PnP ADMM framework from Sureau et al. (2020) with the DRUNet denoiser from Zhang et al. (2022). Specifically, we employ DRUNet in the proximal update step of the PnP framework to enhance the denoising performance.
Algorithm 1 Plug-and-Play ADMM algorithm to deconvolve a galaxy image, inspired by Sureau et al. (2020)
1:  Initialize: Set $\rho_{0}=1,\rho_{max}=10,\eta=0.5,\gamma=1.4,\Delta_{0}=0$, $\mathbf{x}^{(0)}=\mathbf{y},\mathbf{z}^{(0)}=\mathbf{x}^{(0)},\boldsymbol{\mu}^{(0)}=0,\epsilon$
2:  for $k=0$ to $N_{iterations}$ do {Main Loop}
[…]
5:     Lagrange multiplier update: $\boldsymbol{\mu}^{(k+1)}=\boldsymbol{\mu}^{(k)}+\left(\mathbf{x}^{(k+1)}-\mathbf{z}^{(k+1)}\right)$
6:     $\Delta_{k+1}=\frac{1}{\sqrt{n}}\left(||\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}||_{2}+||\mathbf{z}^{(k+1)}-\mathbf{z}^{(k)}||_{2}+||\boldsymbol{\mu}^{(k+1)}-\boldsymbol{\mu}^{(k)}||_{2}\right)$
7:     if $\Delta_{k+1}\geq\eta\Delta_{k}$ and $\rho_{k+1}\leq\rho_{max}$ then
8:        $\rho_{k+1}=\gamma\rho_{k}$
9:     else
10:        $\rho_{k+1}=\rho_{k}$
11:     end if
12:  end for
13:  return  $\left\{\mathbf{x}^{(k+1)}\right\}$
B.1 The proposed solution
Considering the forward model described in Section 2.1, we defined the following loss functions, akin to equations 6-8, but incorporating an additional augmented Lagrangian term with the dual variable. These functions were then minimized using the algorithm outlined in Section B.2.
$L_{r}(\mathbf{x}_{r})$ $=\frac{1}{2}\left\|\frac{\mathbf{h}_{r}\ast\mathbf{x}_{r}-\mathbf{y}_{r}}{\sigma_{r}}\right\|_{F}^{2}+\frac{\rho}{2}\left\|\mathbf{x}_{r}-\mathbf{z}_{r}+\mathbf{\mu}_{r}\right\|_{F}^{2}+\lambda_{r}\textbf{K}$ (17)
$L_{i}(\mathbf{x}_{i})$ $=\frac{1}{2}\left\|\frac{\mathbf{h}_{i}\ast\mathbf{x}_{i}-\mathbf{y}_{i}}{\sigma_{i}}\right\|_{F}^{2}+\frac{\rho}{2}\left\|\mathbf{x}_{i}-\mathbf{z}_{i}+\mathbf{\mu}_{i}\right\|_{F}^{2}+\lambda_{i}\textbf{K}$ (18)
$L_{z}(\mathbf{x}_{z})$ $=\frac{1}{2}\left\|\frac{\mathbf{h}_{z}\ast\mathbf{x}_{z}-\mathbf{y}_{z}}{\sigma_{z}}\right\|_{F}^{2}+\frac{\rho}{2}\left\|\mathbf{x}_{z}-\mathbf{z}_{z}+\mathbf{\mu}_{z}\right\|_{F}^{2}+\lambda_{z}\textbf{K}$ (19)
where $\text{where }\textbf{K}=\left\|\frac{\hskip 2.0pt\mathbf{h}_{euc}\ast\sum\limits_{c\in\{r,i,z\}}\alpha_{c}\mathbf{x}_{c}-\mathbf{y}_{euc}}{\sigma_{euc}}\right\|_{F}^{2}$ (20)
As before, the first terms in equations 17-19 represent the data fidelity terms for each respective band, with $\sigma_{r}$, $\sigma_{i}$, $\sigma_{z}$ being the corresponding noisemaps. The second terms represent the augmented Lagrangian, incorporating the dual variable $\mathbf{z}$ to split the problem into two sub-problems: an inversion/deconvolution step followed by a denoising step. The third terms, as explained in Section 2.2, are the constraining terms that enforce the condition that the sum of images in the Rubin $r$-, $i$-, and $z$-band images equals the $Euclid$ $VIS$-band images. For the ADMM update, the parameter $\rho$ was manually tuned based on the approach from Sureau et al. (2020) to strike a balance between quickly stabilizing the algorithm (with a higher $\rho$) and prioritizing the minimization of the data fidelity term in the early iterations (with a lower $\rho$). The values of all other hyperparameters have been previously described in Sections 2.2 and 2.3.
B.2 The algorithm
The PnP ADMM algorithm inspired by Sureau et al. (2020) is summarized in Table 1. The first step consists of solving the loss functions 17-19 alternatively at iteration $k$ using the accelerated iterative convex algorithm FISTA (Beck & Teboulle 2009). In the second step, the DRUNet denoiser functions as a projector in the proximal update step, as described earlier. The final step regulates the augmented Lagrangian parameter, ensuring its increase when the optimization parameters exhibit insufficient change.
B.3 Results
The algorithm simultaneously processed the noisy simulations from the three Vera C. Rubin bands and the Euclid $VIS$-band, along with their respective PSFs. These noisy images served as initializations or first guesses. The algorithm was run for 200 iterations, with convergence typically observed within $150$-$200$ iterations for all images in our dataset. Figure 11 shows the convergence plot of the loss function for the deconvolved output in Figure LABEL:subfig:mcdec1_pnp.
Figure 11: The loss function for the galaxy shown in Figure LABEL:subfig:mcdec1_pnp. Convergence is achieved at around 200 iterations when the relative change in loss value $<10^{-3}$ and the curve is flat.
Figure LABEL:fig:deconv_outputs_pnp presents the same two examples of deconvolved images as shown in Figure LABEL:fig:deconv_outputs. Qualitatively, the outputs for the two methods closely resemble each other. Compared to the original algorithm, there is a slight reduction in NMSE by approximately $0.26\%$. However, due to the additional computational steps involved in the iterative process and the proximal denoising step, PnP ADMM takes approximately 50 times longer to run than the original algorithm. This is consistent with the findings of Sureau et al. (2020). Hence, this significant increase in computational time makes it less practical for use on large datasets.